Kelly wants to lose 55 pounds over the course of a year (365 days). How many calories does she need to eliminate each day to meet this goal?
Group of answer choices
527 (rounded)
328 (rounded)
489 (rounded)
652 (rounded)
Flag this Question
Question 7 3 pts
Sam eats an extra 650 kcals each day for 180 days. How many pounds can he expect to gain during this time?
Group of answer choices
33.4
67
105.6
210.2
Flag this Question
Question 8 3 pts
Terrence is underweight by 35 pounds. If he eats an extra 750 kcals each day, how long would it take him to gain 35 pounds?
Group of answer choices
145 days (rounded)
163 days (rounded)
211 days (rounded)
189 days (rounded)
Answer #1
  • 527 (rounded) - one pound equal to 3500 calories.

= 55×3500 =1,92,500

=192500÷365

= 527

  • 33.4   
  • 163 days